OpenStudy (anonymous):

what percent of 200 is 161.7? please show all work!

OpenStudy (anonymous):

\[161.7\div200 x 100% = 0.8085 x 100% = 80.85%\]

OpenStudy (anonymous):

161.7 / 200 x 100% = 80.85%

OpenStudy (anonymous):

=(161.7/200)*100=80.85